Many doors in public places are fitted with simple and inexpensive mechanisms that are designed to close the doors while reduM = 70kg C = 1Ns/m k = 26N/m FO = 5N/S 9 = 95 (duration of the couple in seconds) b = 45N (magnitude of the step in N) z = 2.
Answer #1

K O M nose Ex Ax Fkx-cda dt - Mdda dte Given ,M=70 cal 3 Kzg6 F 25 po so their 5-aba-da at toldan ats to dla dta tabr 25 Takeopen loop transfer function x@) G(S) 45(5+2.4) (5 +4-3) [søt +26 6 ] 70 poles = -4.3 S = -0.00714 tj0.6094 3g = -0.00714 -30.(3) centroid = Real part of alpole-sheal partito..zero TZ P. 2 = -4.3 - g 0.00714)-4-204) g =-0.95714 There is no break in or